A helicopter is flying in a straight line over a l
Answer #1

Initial horizontal speed = vx = 5.70 m/s

Height above ground = y = 8.9 m

Initial velocity of package with respect to helicopter = xph = = -14 m

Initial velocity of package with respect to helicopter = Initial velocity of packet with respect to Earth - Speed of helicopter

-14.0 =v - 5.70 m/s

v=- 8.3 m/s

speed = 8.3 m/s

(b) time taken to strike the ground = t = ?

vertical distance travelled = y = 8.90 m

initial vertical velocity = voy = 0 m/s

acceleration = a = -9.81 m/s^2

Using

y = voyt +0.5at^2

-8.9 = 0+ 0.5(-9.81)*t^2

t = 1.81 s

in this time, package hits ground, and helicopter travels d = speed of helicopter x time = 5.70 x 1.81 = 10.3 m

package covers horizontal distance = d' = speed of package x time = 8.3 x 1.81 = 15 m

horizontal distance between package and helicopter when package hits ground = 15 - 10.3 = 4.7 m

(c) angle = arctan( final vertical velocity/ horizontal velocity)

final vertical velocity = v'

initial vertical velocity = 0

acceleration = a - -9.81 m/s^2

t = times = 1.81 s

v' -v = at

v' - 0 = -9.81 x 1.81 = -17.8 m/s

angle = arctan ( -17.8/8.3) = 65 degrees with ground